OpenStudy (anonymous):

The area of a square is found by squaring one of its sides, (A = s 2). A certain square has an area of z 2 + 18z + 81. Factor the trinomial to find a side of the square. side = z + 27 side = z + 18 side = z + 9

OpenStudy (kc_kennylau):

Use the identity \(\large a^2+2ab+b^2=(a+b)^2\) :)

OpenStudy (kc_kennylau):

Use it to factorize \(\large z^2+18z+81\) :)

jhonyy9 (jhonyy9):

@dirtydan667 dont like you cooperate anything

OpenStudy (nurali):

so to factor the trinomal: z^2+18z+81 In this problem 9*9=81 and 9+9=18, so insert 9 as the right hand term of one factor and 9 as the right-hand term of the other factor. (z+9)(z+9) Combine the two common factors of (z+9) by adding the exponents. (z+9)^2

OpenStudy (nurali):

so since it would factor out to be (z+9)^2 this would be equivalent to (z+9)(z+9) so, there for side = z + 9